OpenStudy (anonymous):

what is the normal curve area, to the nearest whole percent, between z=-0.38 and z=+1.5

OpenStudy (hossamhoussien):

\[P (-0.38<Z<1.5) = P (Z<1.5) - P (Z<-0.38)\] \[P (Z<1.5) = 0.9332\] \[P (Z<-0.38) = 0.3520\] \[0.9332 - 0.3520 = 0.5812\]